You are the beneficiary of a trust fund that will start paying you cash flows in five years. The cash flows will be $25,000 per year and will continue for 40 years. If the interest rate is 4% per year, what is the value needed in the trust fund now to fund these cash flows?
Answer #1

Capital scale 

= present value of the first annuity at t + 1 and lasting for n years of deferred annuity 

= annual cash flow/interest rate * (1-1/(1+rate)^n)*1/(1+rate)^t 

= 25000/4%*(1-1/1.04^40)*1/1.04^4 

= 422973.65

answered by: Gavin
Answer #2

Cash flowe = Time lesson = Interest rate $25,000 5 чеass 16 40 чао за = 4% =0.04 .X - Formula of Net present inalie. NRV = In Cash Flowe NPV = (1+r)* Here, j suns from 5 to 40.. 25000 1925 [1:04%40 NPV = 25000 wf = abowo termine = 2500 leting e tentang --- + (1 16 GP - Using, sum of GP :- II- & a 1st term { → common caNPV = $ 25000 / 0.2002 x 3-1039 126 = $ 250VD x 0.6214 x 26 = $ 4,03,910 . So - The value needed in the trust fund now to fun​​​​​​Value needed now is $4,03,910